The trailer is driving 68km/h.The tractor 3 tons is loaded on the trailer. When trailer suddenly stops, how much force apply on the tractor and push forward?
1
Expert's answer
2019-05-17T11:50:45-0400

The impulse produced by force which pushes a tractor forward

"J=F\Delta t=|mv_f-mv_i|"

"=|3000\:\rm{kg}\times 0\:\rm{m/s}-3000\:\rm{kg}\times 0\:\rm{m/s}\times 68\:\rm{km/h}\times \frac{1000\:\rm{m/km}}{3600\:\rm{s/h}}|""=56700\:\rm{N\:s}"

The force value is dependent on braking time

"F=\frac{J}{\Delta t}"